OpenStudy (lgbasallote):

if you have to ask, you'll never know. If you know, you need only ask simple arithmetic: 1 + 1 = 3 2 + 2 = 3 3 + 3 = 3 4 + 4 = 5 5 + 5 = ? show solution

OpenStudy (jlastino):

7?

OpenStudy (lgbasallote):

lol no =))) try again

OpenStudy (experimentx):

5

OpenStudy (lgbasallote):

how?

OpenStudy (experimentx):

1 + 1 = 3 2 + 2 = 3 3 + 3 = 3 <--- three times repeated 4 + 4 = 5 <--- only once??? that's not fair.

OpenStudy (lgbasallote):

lol @asnaseer =)))) @experimentX answer is right but wrong solution haha

OpenStudy (lgbasallote):

haha lol closE :P

OpenStudy (lgbasallote):

it's actually more complex :P 5 + 5 = 5 actually...i'm waiting for the right solution hehe

OpenStudy (jlastino):

hahahaha count the number of vowels when you spell out .

OpenStudy (jlastino):

five plus five = 5 vowels

OpenStudy (lgbasallote):

haha still not the formula :P

OpenStudy (lgbasallote):

not really "word" play.... Hint: 6 + 6 = 3

OpenStudy (jlastino):

this is by far the most challenging brain teaser i've seen this year haha

OpenStudy (experimentx):

because you made it that way ... LOL

OpenStudy (lgbasallote):

lol =)))) there is a formula to this =))

OpenStudy (experimentx):

OIGU

OpenStudy (lgbasallote):

lol not fibonacci =)) i dont even get that haha what is OIGU though?

OpenStudy (experimentx):

ahahah ... that's correct.

OpenStudy (lgbasallote):

ahh lol =)) i thought it was a famous sequence hahaha

OpenStudy (jlastino):

wait! let me think of this until lunch

OpenStudy (experimentx):

It's gonna be famous after I become famous.

OpenStudy (lgbasallote):

@asnaseer i think so..you ALMOST got it a while ago like i said..@Jlastino isnt that a lot of hours away? goodluck with that @experimentX haha

OpenStudy (precal):

is this modern algebr? mod

OpenStudy (lgbasallote):

hmm i dont think so..it's a riddle after all =))

OpenStudy (experimentx):

it's called lgbariddle

OpenStudy (precal):

ok

OpenStudy (lgbasallote):

haha yeahhhh

OpenStudy (lgbasallote):

that pattern is interesting though..might use that when it has been forgotten >:))))

OpenStudy (lgbasallote):

i dunno bout that...guess you'll just laugh at the absurdity :P lol hahaha

OpenStudy (inkyvoyd):

6+6=?

OpenStudy (lgbasallote):

6 + 6 = 3 like i said somewhere in the thread...

OpenStudy (inkyvoyd):

Sorry, missed it

OpenStudy (lgbasallote):

so..do you know the formula? /:)

OpenStudy (inkyvoyd):

In no way. Is it simple arithmetic, and do the two sentences have any relevancy?

OpenStudy (lgbasallote):

nooo..i usually write headings but i cant think of any so i ripped it off harry potter lol and yeah it is simple arithmetic @asnaseer clue: 7 + 7 = 3

OpenStudy (inkyvoyd):

ANyting besides 3 and 5?

OpenStudy (lgbasallote):

nahhh..all numbers up to ten add up to 3 and 5

OpenStudy (saifoo.khan):

11?

OpenStudy (lgbasallote):

lol wrong :P

OpenStudy (inkyvoyd):

modular arith?

OpenStudy (lgbasallote):

lol i dunno what that is but im pretty sure this is just simple arithmetic with a twist >:))

OpenStudy (inkyvoyd):

4mod3=1 6mod2=0 basically take the remainder after division.

OpenStudy (lgbasallote):

since no one was able to answer the problem after 5 hours ill the answer now...

OpenStudy (lgbasallote):

1 + 1 = one + one = 3 letters + 3 letters = 6 = six = 3 letters 2 + 2 = two + two = 3 letters + 3 letters = 6 = six = 3 letters 3+3 =three + three = 5 letters + 5 letters = 10 = ten = 3 letters 4 + 4 = four + four = 4 letters + 4 letters = 8 = eight = 5 letters 5 + 5 = five + five = 4 letters + 4 letters = 8 = eight = 5 letters @asnaseer @experimentX @Jlastino @saifoo.khan @inkyvoyd

OpenStudy (inkyvoyd):

Eh, lol

OpenStudy (inkyvoyd):

Nice riddle. I just remembered why I hate riddles.

OpenStudy (lgbasallote):

hahahaha =))) they are good exericises though :P

OpenStudy (inkyvoyd):

Yea, 3 years in Asia has made me forget the English language. Three years ago, I would've gotten that after some time. Now, I probably wouldn't ever have gotten the riddle.

OpenStudy (lgbasallote):

hahahaha =)))) what's your nationality?

OpenStudy (inkyvoyd):

Well, I grew up in America, but my parents are from Taiwan...

OpenStudy (lgbasallote):

ahh.i guess te rdiddle is disadvantageous for non-english speakers..

OpenStudy (inkyvoyd):

This is an English website, and it is called a riddle in English, so I really just didn't pay attention to what I asked for.

OpenStudy (lgbasallote):

@experimentX

OpenStudy (jlastino):

nice rule xD

OpenStudy (callisto):

@lgbasallote Why did you copy FFM's words....?

OpenStudy (lgbasallote):

ha lol @Callisto we just copied the same text...it's a harry potter quote

OpenStudy (callisto):

I see.. Haven't read Harry Potter yet :P